Project

General

Profile

RunPlan20150328 » History » Revision 12

Revision 11 (Boyd, Steven, 26 March 2015 15:04) → Revision 12/15 (Heidt, Christopher, 29 March 2015 14:13)

h1. Target commissioning + CKOV HV scan - 28th/29th March 2015 

 h2. The purpose of the run is to: 

 * Get the system to a state where ToF (and CKOV) Calibration data can be taken 
 * Take some TOF calibration data  

 h3. Daily schedule 

 | Date             | Saturday     | Saturday            |     Sunday        | Sunday       |   
 | estimated time | 8am - 4pm      | 4 - 12pm          |    8am - 4pm      |    4pm - 12pm | 
 | MOM              | Y. Karadzhov | Y. Karadzhov      |    Y. Karadzhov | Y. Karadzhov     | 
 | Lead shifter     | P. Franchini | P. Hanlet         |    P. Frachini    | P. Hanlet         | 
 | Shifter          | H. Nebrensky | P. Soler          |    H. Nebrensky | P. Soler |  
 | BLOC             |                |                   |                 |                  |  
 | SOC              |    R. Bayes      | R. Bayes          |    R. Bayes       | R. Bayes         | 
 | Expert           |    Y Karadzhov (DAQ)|    | | | 
 |                  |                |                   |                 |                  |  
 | Aim              | Take Data      |                   |                 |                  | 
 |                  |                |                   |                 |                  | 
 | Beam?            | Yes            | Yes               | Yes             | Yes              |  
 | Additional Personnel            |                   |                 |                  |    | 

 h2. Prerequisites 

 "Run Preparation List":http://micewww.pp.rl.ac.uk/projects/operations/wiki/RunPreparationList 

 h2. Data Requested 

 h2. Beamline Configurations 

 * N.B. Decay Solenoid will not be available. 
 * We expect to be running with 2V beamloss and beam bump active. 

 h3. Preparation 

 * We will run beam into the hall so need to secure the MICE Hall 
 * Target needs to be ready 
 * BeamLine elements need to be ready 
 * Detectors need to be ready 
 * Controls needs to be ready 
 * DAQ needs to be ready 
 * OnMon and OnRec need to be ready 
 * Data Mover needs to be ready 

 h3. Shifter target training 

 * at shift handover on Saturday - somewhere between 15:00-17:00 

 h2. DAQ testing 

 * We will take data with TOF0, TOF1 and Ckov. This means that we will use only one LDC (VME crate). 
 * The main goal of this run is the continue with the commissioning of the new trigger system, and to investigate the stability issue found last weekend.  
 * We will start with the debugging of the the stability issue. This operation will require unspecified amount of time. 

 Reference Run without Decay Solenoid 

 | Number of pulses | Particle Species| p at Tgt | p@D1| p@D2| p@Tof0 |Proton Absorber | Q1 | Q2 | Q3 | D1| DS | D2| Q4 | Q5 | Q6 | Q7 | Q8 | Q9 | Similar Run(s)| 
 | # |    | MeV/c |MeV/c |MeV/c | MeV/c | mm | A| A| A| A| A| A| A| A | A | A | A | A| DAQ run #| 
 | 1 000 |    pion| 272.42|268.67|265.98| 143.4| 29 | 57.8 | 105.6 | 64.6 | 202.8| ---| 100.4| 176.8 | 237.1| 157.2| 157.6| 238.5 | 203.7| 5729 | 

 #    The beam line will be first set for Reference Run without Decay Solenoid. 
 # Then we will continue with the Ckov momentum scan on the pion momentum 

 h2. CKOV momentum scan run 

 * We will do a momentum scan on the pion momentum. There should be enough statistics with only 2000 actuations each run. 
 * To be carried out simultaneously with the debugging of the stability issue. 

 | Number of pulses | Particle Species| p at Tgt | p@D1| p@D2|Proton Absorber | Q1 | Q2 | Q3 | D1| DS | D2| Q4 | Q5 | Q6 | Q7 | Q8 | Q9 | Similar Run(s)| 
 | #    |    | MeV/c |MeV/c |MeV/c | mm | A| A| A| A| A| A| A| A | A | A | A | A| DAQ run #| 
 | 2000 |    pion| 300 |268.67|265.98| 29 | 57.8 | 105.6 | 64.6 | 202.8| ---| 100.4| 176.8 | 237.1| 157.2| 157.6| 238.5 | 203.7| 5729 | 
 | 2000 |    pion| 325 |321.5|319| 83 | 69.1 |126.3 |77.31 | 246| ---| 119.6| 212.4| 284.8 | 188.9 | 193.7| 293.2 |250.6| 5439 | 
 | 2000 |    pion| 350 |346.58 | 344.14 | 83 | 74.5 | 136.1 | 83.3| 267.7 | ---| 128.8| 229.2 | 307.4 | 204|    210.6 | 318.8 | 272.5 | 5444 | 
 | 2000 |    pion| 375 |371.6 | 369.2 | 83 | 79.8 | 146 | 89.4 | 290.4 | ---| 137.9| 246 | 330 | 218.9|    227.3 | 344.1 | 294.2 | 5442 | 
 | 2000 |    pion| 400 |396.7 | 394.3 | 83 | 85.2 | 155.8 | 95.4 | 314.3 | ---| 147.1| 262.8 | 352.5 | 233.9|    244 | 369.4 | 315.8 | 5440 | 
 | 2000 |    pion| 425 |421.7 | 419.36 | 83 | 90.55 | 165.6 | 101.38 | 339.36 | ---| 156.26| 279.61 | 374.97 | 248.84| 260.49 | 394.43 | 337.23 | 5437 | 
 | 

 h2. TOF calibration by using the old trigger system. 

 * This calibration will be used for comparison New vs. Old Trigger 
 * The goal will be to take 30 000 target pulse or more. 
 * Trigger provided by the old Tr. system. 
 # The beam line will be set for    TOF Calibration run. 

 h3. TOF Calibration run 

 | Number of pulses| Beamloss |    Particle Species| p at Tgt | p@D1| p@D2| Proton Absorber | Q1 | Q2 | Q3 | D1| DS | D2| Q4 | Q5 | Q6 | Q7 | Q8 | Q9 | Similar Run(s)| 
 | #    | V    | mm | A| A| A| A| A| A| A| A | A | A | A | A|A|A|A|A| DAQ run #| 
 | 30000 | 1 | Positron | 300.0 | 259.7 | 249.58 | 29    | 56.35 | 102.7 | 62.65 | 195.72| ---| 98.59| 163.16 |218.79 | 114.14| 83.40| 125.70 | 80.06| 1678 | 

 h2. -CKOV CKOV HV scan run- run 

 * We also want to do a Ckov HV and momentum scan. 
 * To be carried out simultaneously and parasitically with the TOF calibration run on electron beam as shown above 
 * Previous run to be split into 5 runs with 5000 pulses each run, scanning in a range of +- 50 V around nominal settings for CKOV HV settings 

 | Run type | Number of pulses| CKOVa1 | CKOVa2 | CKOVa3 | CKOVa4 | CKOVb1| CKOVb2 | CKOVb3 | CKOVb4 | 
 |    | #    | V    | V | V | V | V | V | V | V |  
 | Default CKOV HV | 5000 | 1610 | 1520    | 1570 | 1625 | 1540 | 1590 | 1540 | 1495| 
 | Default CKOV HV - 50V | 5000 | 1560 | 1470    | 1520 | 1575 | 1490 | 1540 | 1490 | 1445| 
 | Default CKOV HV - 25V |    5000 | 1585 | 1495    | 1545 | 1600 | 1515 | 1565 | 1515 | 1470| 
 | Default CKOV HV + 25V | 5000 | 1635 | 1545    | 1595 | 1650 | 1565 | 1615 | 1565 | 1520| 
 | Default CKOV HV + 50V | 5000 | 1660 | 1570    | 1620 | 1675 | 1590 | 1640 | 1590 | 1545| 

 h2. Finish 

 * Shutdown target and re-open Hall for Monday